OpenStudy (anonymous):

Linear Algebra Question: Find the polynomial of degree 4 whose graph goes through the points (-3, -227), (-1, -3), (0, 4), (1, 13), and (2, 18). f(x) = ?

myininaya (myininaya):

just plug in and you have your polynomial

OpenStudy (anonymous):

thank you very much :)

myininaya (myininaya):

\[f(x)=f(x_0) \cdot \frac{x-x_1}{x_0-x_1} \cdot \frac{x-x_2}{x_0-x_2} \cdot \frac{x-x_3}{x_0-x_3} \cdot \frac{x-x_4}{x_0-x_4}\] \[+f(x_1) \cdot \frac{x-x_0}{x_1-x_0} \cdot \frac{x-x_2}{x_1-x_2} \cdot \frac{x-x_3}{x_1-x_3} \cdot \frac{x-x_4}{x_1-x_4}\] \[+f(x_2) \cdot \frac{x-x_0}{x_2-x_0} \cdot \frac{x-x_1}{x_2-x_1} \cdot \frac{x-x_3}{x_2-x_3} \cdot \frac{x-x_4}{x_2-x_4}\] \[+f(x_3) \cdot \frac{x-x_0}{x_3-x_0} \cdot \frac{x-x_1}{x_3-x_1} \cdot \frac{x-x_2}{x_3-x_2} \cdot \frac{x-x_4}{x_2-x_4}\] \[+f(x_4) \cdot \frac{x-x_0}{x_4-x_0} \cdot \frac{x-x_1}{x_4-x_1} \cdot \frac{x-x_2}{x_4-x_2} \cdot \frac{x-x_3}{x_4-x_3}\]

OpenStudy (anonymous):

Actually, I don't understand how to plug this in. the answer they are looking for is in the form of __(x^4) + __(x^3) + __(x^2) + __x + __

myininaya (myininaya):

its too long to write on latex are you sure you don't know how to plug in?

myininaya (myininaya):

what is f(x_0)?

OpenStudy (anonymous):

is it -227?

myininaya (myininaya):

yes

myininaya (myininaya):

what is x_1?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

-3

myininaya (myininaya):

x_1=-1

myininaya (myininaya):

f(x_1)=-3

OpenStudy (anonymous):

oh my bad. yes

OpenStudy (anonymous):

f(x_1) is -3. read too fast

myininaya (myininaya):

see you know how to use a formula lol

myininaya (myininaya):

do you know what this is called?

myininaya (myininaya):

its called lagrange polynomial

myininaya (myininaya):

there may be other ways to find the polynomial but i remember lagrange best

OpenStudy (anonymous):

no ive never seen anything so complexly put :P

myininaya (myininaya):

lol

myininaya (myininaya):

what way have you been using?

OpenStudy (anonymous):

i think my teacher wanted it done with augmented matrices

myininaya (myininaya):

i can't remember that way but if you want i can give you the answer using my way (well not my way, but by way lagrange) just give a few minutes

OpenStudy (anonymous):

ill do it too :)

myininaya (myininaya):

i didn't feel like multiplying it all out and combining like terms lol

OpenStudy (anonymous):

:D ty

OpenStudy (zarkon):

Row reduce the matrix \[\left[\begin{matrix}(-3)^4 & (-3)^3 & (-3)^2 & -3 & 1 & -227 \\ (-1)^4 & (-1)^3 & (-1)^2 & -1 & 1 & -3 \\(0)^4 & (0)^3 & (0)^2 & 0 & 1 & 4 \\(1)^4 & (1)^3 & (1)^2 & 1 & 1 & 13 \\(2)^4 & (2)^3 & (2)^2 & 2 & 1 & 18\end{matrix}\right]\]

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

do i consider (-3)^4 = -3 or 81?

OpenStudy (zarkon):

you get \[\left[\begin{matrix}1 & 0 & 0 & 0 & 0 & -2 \\ 0 & 1& 0& 0& 0& 3 \\0& 0 &1 & 0 & 0 & 3 \\0 & 0 & 0 & 1 & 0 & 5 \\0 & 0 & 0 & 0 & 1 & 4\end{matrix}\right]\]

OpenStudy (zarkon):

\[(-3)^4=81\]

OpenStudy (zarkon):

then \[f(x)=-2x^4+3x^3+3x^2+5x+4\]

OpenStudy (anonymous):

wow that was fast :D. im still doing row operationS :P

OpenStudy (anonymous):

tysvm

OpenStudy (zarkon):

you get the same solution if you use the Lagrange Polynomial (as myininaya used)

myininaya (myininaya):

i think i remember that way,zarkon lol

OpenStudy (zarkon):

it is pretty simple... let \[f(x)=a_4x^4+a_3x^3+a_2x^2+a_1x+a_0\] plug in all five of the points given above into the function. you get a system of 5 equations and 5 unknowns...then just solve the system.

myininaya (myininaya):

when i let wolfram put into standard form I'm not getting what you have -227x(x+1)(x-1)(x-2)/[120]-3x(x+3)(x-1)(x-2)/[-12]+4(x+3)(x+1)(x-2)(x-1)/[-6]+13(x+3)(x+1)x(x-2)/[-8]+18x(x+3)(x+1)(x-1)/[30]

OpenStudy (zarkon):

you must have typed something in incorrectly. I have a Lagrange poly program on my calculator and it is giving the same result that i have above

myininaya (myininaya):

so we have (x,y)_0=(-3,-227) (x,y)_1=(-1,-3) (x,y)_2=(0,4) (x,y)_3=(1,13) (x,y)_4=(2,18) ?

myininaya (myininaya):

those are the points i'm using is that what you have?

OpenStudy (zarkon):

yes

myininaya (myininaya):

dang it ...

OpenStudy (zarkon):

i found your error...

myininaya (myininaya):

i'm a horrible typer

myininaya (myininaya):

me too

myininaya (myininaya):

6 instead of -6

OpenStudy (zarkon):

yep

myininaya (myininaya):

ok but anyways both processes should yield the same answer (if you make no mistakes or typo's) hopefully they will let you use a program for something like this that's too much algebra by hand

myininaya (myininaya):

when i was doing this stuff they made us make our own programs like in maple or something like that

myininaya (myininaya):

but you have to understand what the process is to make the program

OpenStudy (zarkon):

I used matlab..when I was taking numerical analysis

OpenStudy (zarkon):

yep

myininaya (myininaya):

i prefer maple

OpenStudy (zarkon):

I don't like to program in maple

myininaya (myininaya):

you like to program in your nspire

OpenStudy (zarkon):

matlab is so easy to make programs with

OpenStudy (zarkon):

yes I do

OpenStudy (zarkon):

it has a nice programming language

myininaya (myininaya):

i don't have enough money i will wait on the nspire :(

OpenStudy (zarkon):

I got it as an x-mas gift

myininaya (myininaya):

that is why christmas is great because of all the presents

OpenStudy (zarkon):

my mother-in-law spends way too much money on me during Christmas :)

myininaya (myininaya):

someone let me borrow their n-spire and i accidently dropped just once and i broke it so i have to buy them an n-spire lol i will never borrow anything ever again from anyone

OpenStudy (zarkon):

lol...that's not good

myininaya (myininaya):

i had it for one freakin' day

OpenStudy (zarkon):

crazy.

myininaya (myininaya):

one drop=no work wth

OpenStudy (zarkon):

yeah...it's not like the 83...you drop the 83 and it will be fine

myininaya (myininaya):

my ti83 i dropped that numerous times stills works but still stolen

OpenStudy (zarkon):

yep

myininaya (myininaya):

so i'm getting a ti83 again no upgrade

OpenStudy (zarkon):

you should get an 83+...it has flash memory

myininaya (myininaya):

actually i think it might have been plus

myininaya (myininaya):

we will see what i get in the mail in a few days

myininaya (myininaya):

i could go to my ebay account and look it up but i would rather be surprised :)

OpenStudy (zarkon):

lol...nice

myininaya (myininaya):

maybe they will put some really cool calculator program on the ipad one day

OpenStudy (zarkon):

probably

myininaya (myininaya):

that would be so cool to take your finger and point on the screen where you want the graph shaded or something like that

OpenStudy (zarkon):

that would be nice

myininaya (myininaya):

so algebra students won't have to think what is the inequality here: